OpenStudy (anonymous):

If m varies directly as the square root of y, inversely as p^2, and directly as n, what happens to m when y is quadrupled, p is tripled, and n is tripled? So this is what I got so far: m= (k*n*sqrt of y)/p^2

OpenStudy (tkhunny):

It is an algebra problem. \(m = k\dfrac{(3\cdot n)\sqrt{4\cdot y}}{(3\cdot p)^{2}} = \left(k\cdot\dfrac{n\sqrt{y}}{p^{2}}\right)\cdot What? = m\cdot What?\)

OpenStudy (anonymous):

Yeah it's Algebra 2 and all variation equations. I understood every problem I had to do except this one. So would it be multiplied by p^2 to both sides?

OpenStudy (tkhunny):

?? No, just resolve all the new constants.

OpenStudy (anonymous):

Sorry I got confused. Thank you!